# Geometry of quantum states and chaos-integrability transition

### v1

@@ Abstract
We consider the geometry of quantum states associated with classes of random matrix Hamiltonians, in particular ensembles that show integrability to chaotic transition in terms of the nearest neighbour energy level spacing distribution. In the case that the total Hamiltonian contains a single parameter, the distance between two states is captured by the fidelity susceptibility, whereas, when the total Hamiltonian contains multiple parameters, this distance is given by the quantum metric tensor. Since the fieldity susceptibility is closely related to the two-point correlation function, we first calculate the relevant correlation functions of a random matrix belonging to the Gaussian unitary ensemble in terms of the spectral form factor of the total Hamiltonian, show how to obtain the fidelity susceptibility from this correlation function, and explain the role played by energy level correlation. Next, by performing suitable coordinate transformations, we solve the geodesic equations corresponding to the quantum metric tensor obtained from an integrability-breaking random matrix Hamiltonian and obtain the geodesic distance between two points on the parameter manifold to show that any point far away from the integrable phase can be reached by a finite value of this distance. Finally, we obtain and discuss different properties of the fidelity susceptibility associated with Hamiltonians belonging to another random matrix ensemble which shows integrability to chaos transition, namely the Gaussian $\beta$-ensembles with general values of the Dyson index $\beta$, and show that the fidelity susceptibility shares generic features with the first class of Hamiltonians.
@@ I Introduction
In the context of quantum phase transitions, the QMT has emerged as a sensitive probe of criticality, capturing singular behaviour near critical points even when conventional order parameters are absent [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Meanwhile, the Berry curvature part of the QGT encodes the system’s response to adiabatic perturbations and underlies a variety of topological invariants, such as Chern numbers, which, in the linear response theory formalism, can be attributed to the QGT contributions to the generalized susceptibilities and dynamical response functions, linking geometric properties of quantum states to measurable quantities [19, 20, 21]. In this paper, we use the geometry of quantum states of a random matrix Hamiltonian belonging to an ensemble of matrices to characterise the transition from integrability to chaos in these ensembles.
@@
However, in many physical situations, quantum systems can consist of parts of different symmetries, or can have classically chaotic and integrable counterparts. If a parameter exists in such systems, such that changing it causes the system to make a transition from different symmetry classes, then these transitions would be reflected in the symmetry class of the corresponding random matrix. Therefore, in those cases, the quantum systems can also follow other level spacing statistics, which are intermediate between the Wigner-Dyson statistics of classical Gaussian ensembles and Poisson statistics of integrable systems [25, 26, 27, 28, 29, 30].
@@
In this paper, we use both of these generalised random matrix ensembles to analyse the geometry of quantum states of Hamiltonians drawn from these. In the first part of the paper, after briefly reviewing the definitions of the QGT, QMT and ensemble-averaged QMT in section II, we study the geometric quantities associated with the ensemble-averaged QMT (originally derived in [42]) obtained from a GUE Hamiltonian and an integrability-breaking Hamiltonian that is a generalisation of the RP model. Specifically, in section III and V we find out suitable coordinate transformations which make the QMTs derived in [42] much simpler to work with. In terms of these coordinates, one can then solve the geodesic equations associated with the metric tensor and obtain the geodesic distance between two points on the space of parameters, which helps to understand the chaos-to-integrability transition in terms of this geometric quantity. In particular, we show that any point far away from the integrable point can be reached by a finite geodesic distance.
@@
In the second part, we obtain the geometry of eigenstates belonging to the tridiagonal Gaussian $\beta$-ensembles, which also show integrability to chaos transition with varying Dyson index. Assuming the total Hamiltonian is a sum of two terms, a $2\times 2$ matrix belonging to this ensemble, which is added to a diagonal Hamiltonian, we obtain an analytical expression for the FS for any generic value of the Dyson index, and study its behaviour with respect to the change of the relevant parameters (section VI). We show that the FS shares generic features with those of the generalised RP model. Furthermore, to complement the analysis of the tridiagonal ensembles which break the rotational invariance of the classical Gaussian ensembles, an invariant $2\times 2$ random matrix ensemble with an effective Dyson index is also analysed, and we show that the FS still captures the essential signature of the chaos-to-integrability transition (see section VII).
This paper also contains several appendices where we provide some details that are left out of the main text. Specifically, in Appendix B we have outlined the steps of the derivation of the components of the QGT for the GUE Hamiltonian, and in Appendix D we have provided an alternative derivation of the correlation function computed in section IV, which is directly related to the components of the metric tensor. Finally, in Appendix E, the expression for the FS of a specific random matrix ensemble, whose eigenvalues are drawn from a Gaussian $\beta$-ensemble with generic $\beta$ and has Haar random eigenstates, is obtained, and its properties are discussed.
@@ II Fidelity susceptibility and Quantum metric tensor for random matrix ensembles
To begin with, we consider the simplest scenario, i.e., a system whose total Hamiltonian is given by [^2] We assume that the Hamiltonian $H$ has no degeneracy in the spectrum.
@@
where $r$ is a positive parameter. Now consider the change in the fidelity of the $n$th eigenstate |$\ket{n(r)}$ of the Hamiltonian $H$ in (1), i.e., change in $\mathcal{F}(r,dr)=|\braket{n(r)}{n(r+\delta r)}|^{2}$. [^3] Some authors define the fidelity as the absolute value of the overlap between the eigenstate at a parameter value $r$, and at $r+\delta r$. Expanding the fidelity for small values of $\delta r$, and using the normalisation of the states |$\ket{n(r)}$ one has in the limit $\delta r\rightarrow 0$, the following expansion
@@
$\begin{split}g_{rr}^{n}(r)=\braket{\partial_{r}n(r)}{\partial_{r}n(r)}-\braket{\partial_{r}n(r)}{n(r)}\braket{n(r)}{\partial_{r}n(r)}~{}~{}\\ =\sum_{j(\neq n)}\frac{\mathcal{H}_{nj}\mathcal{H}_{jn}}{(E_{j}(r)-E_{n}(r))^{2}}~{}.~{}\end{split}$ (3)
Here $\mathcal{H}_{nj}=\braket{n}{\mathcal{H}}{j}$ denotes the matrix elements of $\mathcal{H}$ in the basis of |$\ket{n}$, the eigenstate of $H$. Note that this quantity is invariant under an overall constant scaling transformation of the Hamiltonian $H(r)$.
@@
Now assume that $H_{0}$ and $\mathcal{H}$ are independent random matrices of rank $N$, and the probability distribution for the total Hamiltonian ($H$) is given by
@@
Here $P(H_{0})$ and $P(\mathcal{H})$ are the distributions for the independent random matrices $H_{0}$ and $\mathcal{H}$, respectively. In this paper, we mostly consider Gaussian random matrices, for which the potential function is $V(H)=H^{2}$.
@@ II.1 Connection with correlation functions
$g_{rr}^{n}(r)=\frac{1}{2\pi}\int_{-\infty}^{\infty}~{}\frac{\text{d}\omega}{\omega^{2}}~{}\mathcal{S}^{n}(\omega)~{},$ (9)
where
$\begin{split}\mathcal{S}^{n}(\omega)=\int_{-\infty}^{\infty}~{}\text{d}t~{}e^{i\omega t}~{}\Big{(}\braket{n}{\mathcal{H}(t)\mathcal{H}}{n}-|\braket{n}{\mathcal{H}}{n}|^{2}\Big{)}~{},\\ ~{}~{}\text{and}~{}~{}\mathcal{H}(t)=e^{iHt}\mathcal{H}e^{-iHt}~{}.~{}~{}~{}~{}~{}~{}~{}~{}\end{split}$ (10)
Now summing over all the eigenstates |$\ket{n}$ and averaging over the random matrix ensemble, in the case $H(r)$ belongs to one, we have the following relation
$\bar{g}_{rr}(r)=\frac{1}{2\pi}\int_{-\infty}^{\infty}~{}\frac{\text{d}\omega}{\omega^{2}}~{}\mathcal{S}(\omega)~{},$ (11)
where we have defined the spectral function
$\mathcal{S}(\omega)=\int_{-\infty}^{\infty}~{}\text{d}t~{}e^{i\omega t}~{}\Big{(}\braket{\mathcal{H}(t)\mathcal{H}}-\frac{1}{N}\mathbb{E}\sum_{n}|\braket{n}{\mathcal{H}}{n}|^{2}\Big{)}~{},$ (12)
with the symbol $\mathbb{E}$ denoting an ensemble average over the probability distribution of the Hamiltonian $H(r)$, and we have used the following notation to denote the map which combines the trace and the ensemble average,
@@
We continue to use this notation in the following. This map can be thought of as the inner product with respect to the infinite temperature thermofield double state.
@@
and use it in the eq. (11) along with (12). In section IV, we obtain an analytical expression for this correlation function when $H(r)$ belongs to a GUE, and discuss its connection with the SFF, and subsequently show how to obtain the FS from this correlation function.
@@
$\mathcal{S}(\omega)=\frac{2\pi}{N}\mathbb{E}~{}\sum_{n\neq j}|\mathcal{H}_{nj}|^{2}~{}\delta\big{(}E_{j}-E_{n}-\omega\big{)}~{}.$ (15)
@@
$m_{k}(r)=\mathbb{E}~{}\int~{}\text{d}\omega~{}\omega^{k}~{}\mathcal{S}(\omega)=\frac{2\pi}{N}~{}\mathbb{E}~{}\sum_{n\neq j}|\mathcal{H}_{nj}|^{2}~{}\big{(}E_{j}(r)-E_{n}(r)\big{)}^{k}~{},$ (16)
we see that these have direct physical meanings. For example, $m_{-2}(r)$ is nothing but the averaged FS (with a proportionality constant $2\pi$). We also note that the function $\mathcal{S}(\omega)$ is the frequency space expression of the connected correlation function, and is related to the dynamical susceptibility of the system through the Kubo linear response formula.
@@ II.2 Quantum metric tensor
The expressions for the FS discussed so far can be generalised when the system Hamiltonian depends on more than one parameter $r^{1},\cdots,r^{k}$, which we collectively denote as $\mathbf{r}$. In that case, it is convenient to define the following ‘line element’, which incorporates changes in one or more of the parameters,
$ds^{2}\equiv 1-|\braket{n(\mathbf{r})}{n(\mathbf{r}+\text{d}\mathbf{r})}|^{2}\equiv g_{ab}^{n}(\mathbf{r})~{}\text{d}r^{a}\text{d}r^{b}+\mathcal{O}(|\text{d}\mathbf{r}|^{3})~{},$ (17)
where the expression for the QMT for the $n$-th eigenstate can be written as
$g^{n}_{ab}(\mathbf{r})=\sum_{m(\neq n)}\frac{\braket{n(\mathbf{r})}{\partial_{a}H(\mathbf{r})}{m(\mathbf{r})}\braket{m(\mathbf{r})}{\partial_{b}H(\mathbf{r})}{n(\mathbf{r})}}{\big{(}E_{m}-E_{n})^{2}}~{},$ (18)
where we have denoted $\partial_{a}=\partial_{r^{a}}$. For a Hermitian Hamiltonian, as expected, this quantity is real and symmetric in the indices $a$ and $b$.
A related quantity studied in [42] for Hamiltonians belonging to random matrix ensembles is called the averaged QMT, and is defined as [^6] In Appendix. A we have provided the generic expression for the QGT of a state which is a superposition of the eigenstates of the system Hamiltonian.
@@
When the Hamiltonian $H$ belongs to some random matrix ensemble, we perform an average of the QMT components over the ensemble, and define an ensemble-averaged QMT as
@@
Before moving on, a comment on the difference between the quantity defined in (3) and a similar quantity studied in the context of random matrices in [49] is in order. In defining the quantity (3) we have assume $H$ to be total Hamiltonian and consider applying a perturbation term of the form $\delta r~{}\mathcal{H}$ which give rise to the FS in (3) and the expectation value of $\mathcal{H}$ are with respect to the eigenstates |$\ket{n}$ of the total Hamiltonian. On the other hand, one can consider a scenario where $H_{0}$ is the initial Hamiltonian and the term $r\mathcal{H}$ acts as a perturbation. In that case, one can derive an identical-looking expression for the associated FS where the matrix elements of $\mathcal{H}$ are with respect to |$\ket{n_{0}}$ the eigenstates of $H_{0}$, and in the denominator the energy difference term would be $(E^{0}_{n}-E^{0}_{j})^{2}$.
@@ III.1 Components of the quantum metric tensor
We first consider the scenario where the total Hamiltonian belongs to one of the classical Gaussian ensembles, which we take to be GUE here. Specifically, we parametrise the total Hamiltonian of the system such that it can be written as [50, 51]
@@
~$\tilde{\mathcal{H}}_{1}=\cos\phi~{}\mathcal{H}_{1}+\sin\phi~{}\mathcal{H}_{2}~{},~{}~{}\tilde{\mathcal{H}}_{2}=-\sin\phi~{}\mathcal{H}_{1}+\cos\phi~{}\mathcal{H}_{2}~{},$ (23)
and $\mathbf{r}=\{r,\phi\}$, while $E_{n}(\mathbf{r})$ denotes the $n$-th energy of the total Hamiltonian. Note that the expression for the total Hamiltonian becomes $H=H_{0}+r\tilde{\mathcal{H}}_{1}$, and the new variables ~$\tilde{\mathcal{H}}_{i}$ are GUE matrices having the same variance $\sigma^{2}=1/N$.

### v2

@@ Abstract
We consider the geometry of quantum states associated with different classes of random matrix Hamiltonians, in particular ensembles that show integrability to chaotic transition in terms of the nearest neighbour energy level spacing distribution. In the case that the total Hamiltonian contains a single parameter, the distance between two states is captured by the fidelity susceptibility, whereas, when the total Hamiltonian contains multiple parameters, this distance is given in terms of the quantum metric tensor. Since the fidelity susceptibility is closely related to the two-point correlation function, we first calculate the relevant correlation functions of a random matrix belonging to the Gaussian unitary ensemble in terms of the spectral form factor of the total Hamiltonian, show how to obtain the fidelity susceptibility from this correlation function, and explain the role played by energy level correlation. Next, by performing suitable coordinate transformations, we solve the geodesic equations corresponding to the quantum metric tensor obtained from an integrability-breaking random matrix Hamiltonian and obtain the geodesic distance between two points on the parameter manifold to show that any point far away from the integrable phase can be reached by a finite value of this distance. Finally, we obtain and discuss different properties of the fidelity susceptibility associated with Hamiltonians belonging to another random matrix ensemble which shows integrability to chaos transition, namely the Gaussian $\beta$-ensembles with general values of the Dyson index $\beta$, and show that the fidelity susceptibility shares generic features with the first class of Hamiltonians.
@@ I Introduction
In the context of quantum phase transitions, the QMT has emerged as a sensitive probe of criticality, capturing singular behaviour near critical points even when conventional order parameters are absent [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Meanwhile, the Berry curvature part of the QGT encodes the system’s response to adiabatic perturbations and underlies a variety of topological invariants, such as Chern numbers, which, in the linear response theory formalism, can be attributed to the QGT contributions to the generalized susceptibilities and dynamical response functions, linking geometric properties of quantum states to measurable quantities [19, 20, 21, 22]. In this paper, we use the geometry of quantum states of a random matrix Hamiltonian belonging to an ensemble of matrices to characterise the transition from integrability to chaos in these ensembles.
@@
However, in many physical situations, quantum systems can consist of parts of different symmetries, or can have classically chaotic and integrable counterparts. If a parameter exists in such systems, such that changing it causes the system to make a transition from different symmetry classes, then these transitions would be reflected in the symmetry class of the corresponding random matrix. Therefore, in those cases, the quantum systems can also follow other level spacing statistics, which are intermediate between the Wigner-Dyson statistics of classical Gaussian ensembles and Poisson statistics of integrable systems [26, 27, 28, 29, 30, 31, 32].
@@
In this paper, we use both of these generalised random matrix ensembles to analyse the geometry of quantum states of Hamiltonians drawn from these. In the first part of the paper, after briefly reviewing the definitions of the QGT, QMT and ensemble-averaged QMT in section II, we study the geometric quantities associated with the ensemble-averaged QMT (originally derived in [44]) obtained from a GUE Hamiltonian and an integrability-breaking Hamiltonian that is a generalisation of the RP model. Specifically, in section III and V, we find suitable coordinate transformations that make the QMTs derived in [44] much simpler to work with. In terms of these coordinates, one can then solve the geodesic equations associated with the metric tensor and obtain the geodesic distance between two points on the space of parameters, which helps to understand the chaos-to-integrability transition in terms of this geometric quantity. In particular, we show that any point far away from the integrable point can be reached by a finite geodesic distance.
@@
In the second part, we obtain the geometry of eigenstates belonging to the tridiagonal Gaussian $\beta$-ensembles, which also show integrability to chaos transition with varying Dyson index. Assuming the total Hamiltonian is a sum of two terms, a $2\times 2$ matrix belonging to this ensemble, which is added to a diagonal Hamiltonian, we obtain an analytical expression for the FS for any generic value of the Dyson index, and study its behaviour with respect to the change of the relevant parameters (section VI). We show that the FS shares generic features with those of the generalised RP model. Furthermore, to complement the analysis of the tridiagonal ensembles which break the rotational invariance of the classical Gaussian ensembles, an invariant $2\times 2$ random matrix ensemble with an effective Dyson index is also analysed, and we show that the FS still captures the essential signature of the chaos-to-integrability transition (see section VII). Finally, in section VIII we analyse the sensitivity of the eigenvalues, rather than the eigenstates, of these systems, by studying the ensemble-averaged curvature of the energy eigenvalues (which is the second derivative of the energy levels with respect to the parameters in the Hamiltonian). Indeed, the sensitivity of the energy levels of a complex many-body quantum system with respect to the parameters is one of the most important approaches developed in the past decades towards an understanding of why energy levels of chaotic quantum systems exhibit random matrix universality [49, 50, 51, 52]. Our main findings and conclusions are summarised in section IX.
This paper also contains several appendices where we provide some important details that are left out of the main text. Specifically, in Appendix B we have outlined the steps of the derivation of the components of the QGT for the GUE Hamiltonian, and in Appendix D we have provided an alternative derivation of the correlation function computed in section IV, which is directly related to the components of the metric tensor. Finally, in Appendix E, the expression for the FS of a specific random matrix ensemble, whose eigenvalues are drawn from a Gaussian $\beta$-ensemble with generic $\beta$ and has Haar random eigenstates, is obtained, and its properties are discussed.
@@ II Fidelity susceptibility and Quantum metric tensor for random matrix ensembles
To begin with, we consider the following generic scenario, where the total Hamiltonian of a quantum system is given by [^2] We assume that the Hamiltonian $H$ has no degeneracy in the spectrum.
@@
[…]
@@
[…]
@@
Now we assume that $H_{0}$ and $\mathcal{H}$ are independent random matrices of rank $N$, and the probability distribution for the total Hamiltonian ($H$) is given by
@@
Here $P(H_{0})$ and $P(\mathcal{H})$ are the distributions for the independent random matrices $H_{0}$ and $\mathcal{H}$, respectively. In this paper, we mostly consider Gaussian random matrices, for which the potential function is $V(H)=H^{2}$. [^6] For a discussion on the distribution of the FS for the Gaussian random matrix ensembles and its comparison with that of the disorder many-body systems, see [55, 56].
@@ II.1 Connection with correlation functions
$g_{rr}^{n}(r)=\int_{-\infty}^{\infty}~\frac{\text{d}\omega}{\omega^{2}}~\mathcal{S}_{n}(\omega)~,$ (9)
where [^7] The standard definition of this function usually does not have the $1/2\pi$ factor. Here we have modified the standard definition so that the subsequent formulas for the FS and the moments are free from the $2\pi$ factor.
[…]
with $\mathcal{H}(t)=e^{iHt}\mathcal{H}e^{-iHt}$. This function has the following spectral representation,
$\mathcal{S}_{n}(\omega)=\sum_{j(\neq n)}|\mathcal{H}_{nj}|^{2}~\delta\big{(}E_{j}-E_{n}-\omega\big{)}~,$ (11)
and has the following moments
$M^{(n)}_{k}(r)=\int~\text{d}\omega~\omega^{k}~\mathcal{S}_{n}(\omega)=\sum_{j(\neq n)}|\mathcal{H}_{nj}|^{2}~\big{(}E_{j}(r)-E_{n}(r)\big{)}^{k}~,~~k\in\mathbb{Z}~.$ (12)
These moments have physical meaning: e.g., for k=2k=-2, the moment $M_{-2}$ is just the FS of the $n$-th eigenstate [47]. The meaning of the moment with k=1k=-1 is discussed in sec. VIII.
[…]
$\bar{g}_{rr}(r)=\int_{-\infty}^{\infty}~\frac{\text{d}\omega}{\omega^{2}}~\mathcal{S}(\omega)~,$ (13)
where we have defined the spectral function (in an appropriate ensemble-averaged sense)
$\mathcal{S}(\omega)=\frac{1}{2\pi}\int_{-\infty}^{\infty}~\text{d}t~e^{i\omega t}~\Big{(}\braket{\mathcal{H}(t)\mathcal{H}}-\frac{1}{N}\mathbb{E}\sum_{n}|\braket{n|\mathcal{H}|n}|^{2}\Big{)}~,$ (14)
with the symbol $\mathbb{E}$ denoting an ensemble average over the probability distribution of the Hamiltonian $H(r)$; and we have used the following notation to denote the map which combines the trace and the ensemble average,
@@
We continue to use this notation in the following. This map can be thought of as the inner product with respect to the infinite temperature thermofield double state, and is also known as the normalised expected trace in the mathematics literature.
@@
and use it in the eq. (13) along with (14). In section IV, we obtain an analytical expression for this correlation function when $H(r)$ belongs to the GUE, and discuss its connection with the SFF, and subsequently show how to obtain the FS from this correlation function.
@@
$\mathcal{S}(\omega)=\frac{1}{N}\mathbb{E}~\sum_{n\neq j}|\mathcal{H}_{nj}|^{2}~\delta\big{(}E_{j}-E_{n}-\omega\big{)}~.$ (17)
@@
$M_{k}(r)=\int~\text{d}\omega~\omega^{k}~\mathcal{S}(\omega)=\frac{1}{N}~\mathbb{E}~\sum_{n\neq j}|\mathcal{H}_{nj}|^{2}~\big{(}E_{j}(r)-E_{n}(r)\big{)}^{k}~,$ (18)
$k$ being integers, we see that these also have direct physical meanings. For example, $M_{-2}(r)$ is nothing but the ensemble-averaged mean FS. We also note that the function $\mathcal{S}(\omega)$ is the frequency space expression of the connected correlation function, and is related to the dynamical susceptibility of the system through the Kubo linear response formula.
@@ II.2 Quantum metric tensor
[…]
$ds^{2}=g_{ab}^{n}(\mathbf{r})~\text{d}r^{a}\text{d}r^{b}~,$ (20)
where $g_{ab}^{n}(\mathbf{r})$, known as the QMT, is the real and symmetric part of the QGT of the $n$th eigenstate, defined as [6, 9]
$\begin{split}G^{n}_{ab}(\mathbf{r})=\braket{\partial_{a}n(\mathbf{r})|\partial_{b}n(\mathbf{r})}-\braket{\partial_{a}n(\mathbf{r})|n(\mathbf{r})}\braket{n(\mathbf{r})|\partial_{b}n(\mathbf{r})}\\ =\sum_{m(\neq n)}\frac{\braket{n(\mathbf{r})|\partial_{a}H(\mathbf{r})|m(\mathbf{r})}\braket{m(\mathbf{r})|\partial_{b}H(\mathbf{r})|n(\mathbf{r})}}{\big{(}E_{m}(\mathbf{r})-E_{n}(\mathbf{r}))^{2}}~.\end{split}$ (21)
The last expression above can be obtained by using the eigenvalue equation for the $n$th eigenstate.
A related quantity studied in [44] for Hamiltonians belonging to random matrix ensembles is called the averaged QGT, and is defined as [^8] In Appendix. A, we have provided the generic expression for the QGT of a state which is a superposition of the eigenstates of the system Hamiltonian.
@@
For a Hermitian Hamiltonian, this quantity is real and symmetric in the indices $a$ and $b$ (its imaginary and antisymmetric part is zero). We call this the mean QMT; however, in the following, we shall often omit the quantifier ‘mean’ and refer to it only as QMT for convenience.
When the Hamiltonian $H$ belongs to some random matrix ensemble, we are interested in the average of the QMT components over the ensemble, so that we define the ensemble-averaged QMT components as
$\bar{g}_{ab}(\mathbf{r})=\mathbb{E}\big{(}g_{ab}(\mathbf{r})\big{)}=\int dH(\mathbf{r})~P(H(\mathbf{r}))~g_{ab}(\mathbf{r})~,$ (23)
@@ III.1 Components of the quantum metric tensor
We first consider the scenario where the total Hamiltonian belongs to one of the classical Gaussian ensembles, which we take to be GUE here. Specifically, we parametrise the total Hamiltonian of the system such that it can be written as [58, 59, 60, 51, 52]
@@
$\tilde{\mathcal{H}}_{1}=\cos\phi~\mathcal{H}_{1}+\sin\phi~\mathcal{H}_{2}~,~~\tilde{\mathcal{H}}_{2}=-\sin\phi~\mathcal{H}_{1}+\cos\phi~\mathcal{H}_{2}~,$ (26)
and $\mathbf{r}=\{r,\phi\}$, while $E_{n}(\mathbf{r})$ denotes the $n$-th energy of the total Hamiltonian. Note that the expression for the total Hamiltonian becomes $H=H_{0}+r\tilde{\mathcal{H}}_{1}$, and the new variables $\tilde{\mathcal{H}}_{i}$ are GUE matrices having the same variance $\sigma^{2}=1/N$.